solve 1
Problem 3: Solve the given problem. 1.7 + 3y + 2y = 2. y + y = e + x, y(0) = 2, 7(0) = 0
Answer #1

yllt zylray = x2 Auxillary equ (02+3D+2)y=0 Let Dam +3m2=0 mi? +2m+m+2=0 m (m+2)+1 (m12)=0 (mutz) (m+I) 20 me-2-1 Yc = q e 2x. -эм ) - Г-н) - 2x +2 2 x | Niw + + 93m 2 3 + ЧР У = Че** + * + ** - 3 + 2 Ч